# User-Feedback-Driven Continual Adaptation for Vision-and-Language Navigation

### v1

@@ III-C Stage 2: User-feedback-driven Trajectory Correction
After executing an instruction $I$, the agent produces a trajectory
@@
ending at viewpoint $v_{T}$. The user then provides feedback: if $v_{T}$ correctly matches the target, the trajectory is accepted; otherwise, the correct goal viewpoint $v^{*}$ is specified.
As illustrated in Stage 2 of Fig. 3, when feedback indicates an error, the corrected path is computed using the A* algorithm on the existing topological graph:
$\tau^{+}=\text{A}^{*}(v_{1}\rightarrow v^{*},G_{E}).$ (9)
Each corrected pair $(I,\tau^{+})$ is stored as a new adaptation sample. To ensure data quality, we remove cases where $\tau^{+}$ is too short (<5<5 nodes) or too long (>7>7 nodes), as these often indicate incomplete topology or inconsistent feedback.
This design transforms minimal user supervision—simply specifying the final viewpoint—into high-quality, in-distribution training data. Compared with unsupervised GSA-VLN adaptation, the resulting dataset provides explicit semantic alignment with user intent, effectively bridging the gap between synthetic instructions and real-world interaction data.
III-D Stage 3: Feedback-based Adaptation and Warm-started Redeployment
@@ III-D1 Memory-bank Warm Start
After fine-tuning, the agent is redeployed in the same environment. Instead of reconstructing the graph and visual features from scratch, it reloads the previously stored memory $\mathcal{M}_{E}=(G_{E},\mathcal{C}_{E},\mathcal{S}_{E})$, enabling immediate reuse of environment topology and cached representations. This reloadable memory extends the GR-DUET design by introducing persistent scene states that can be saved and restored across sessions, maintaining structural and perceptual continuity during redeployment.
@@ III-D3 Hybrid Adaptation
In practical multi-user deployments, several users may interact with the same agent concurrently in one environment. To handle such concurrent supervision efficiently, we introduce hybrid adaptation, which aggregates feedback from all users within a deployment window:
@@
where $\mathcal{D}_{u}$ represents feedback-corrected trajectories from user $u$. By jointly optimizing across multiple user datasets, hybrid adaptation captures diverse linguistic expressions and navigation styles while avoiding bias toward early interactions. Together, continual and hybrid adaptation mechanisms enable the agent to evolve both temporally and socially in long-term, multi-user environments.
By integrating user-feedback-driven dataset construction, imitation-based policy adaptation, and a reloadable memory bank derived from GR-DUET, our framework transforms deployment experience into a sustainable learning loop. The agent continuously improves its navigation ability while efficiently reusing previously explored structural and visual knowledge—achieving adaptive, stable performance beyond static GSA-VLN paradigms.
@@ IV-A Experimental Setup
We conduct experiments on the GSA-R2R dataset [Hong_et_al_2025], which extends the standard Room-to-Room (R2R) benchmark [Anderson_2018] to support scenario adaptation. In this work, we focus on the residential environments and evaluate on two instruction styles: Basic and User. This setting captures the most common real-world deployment scenarios, where navigation agents repeatedly operate in household layouts and encounter both neutral and user-specific language patterns.
@@
We adopt GR-DUET [Hong_et_al_2025] as our primary baseline, since it represents the state of the art for GSA-VLN by explicitly maintaining a global navigation graph for adaptation. To further validate the generality of our approach, we also test other representative VLN methods under the GSA-VLN protocol, including DUET [Chen_DUET], HAMT  [Chen_HAMT_2021], and NaviLLM [Zheng_2024_CVPR]. These methods were originally developed for the standard zero-shot VLN setting, and their performance in the adaptation scenario serves as an additional reference.
@@
All adaptation experiments share consistent optimization and hardware settings. Each adaptation stage consists of 30,000 episodes with a batch size of 2, corresponding to approximately 60,000 optimization steps. Training is conducted on a single NVIDIA RTX 4090 GPU for about 5 days per stage.
@@ IV-B Evaluation Metrics
We evaluate navigation performance using standard metrics from the Vision-and-Language Navigation (VLN) literature, covering goal achievement, path efficiency, trajectory fidelity, and semantic alignment. Following the Room-to-Room (R2R) benchmark [Anderson_2018], we report nine widely used metrics: Success Rate (SR) [Anderson_2018], Oracle Success Rate (Oracle SR)[Anderson_2018], Success weighted by Path Length (SPL) [anderson2018evaluation], Navigation Error (NE) [Anderson_2018], Oracle Error / Oracle Navigation Error [jain2019stay], Path Length [Anderson_2018], Normalized Dynamic Time Warping (nDTW) [ilharco2019general], Success weighted by nDTW (SDTW) [ilharco2019general], and Coverage weighted by Length Score (CLS) [jain2019stay]. These metrics are adopted consistently across recent VLN benchmarks [Anderson_2018, Ku_2020, jain2019stay, Hong_et_al_2025] to ensure comparability with prior work.
@@ IV-C Main Results on the GSA-Basic Test Set
Table I presents the quantitative comparison of different VLN methods on the GSA-Basic test split. We include representative baselines HAMT [Chen_HAMT_2021], NaviLLM [Zheng_2024_CVPR], DUET [Chen_DUET], and GR-DUET [Hong_et_al_2025], where GR-DUET serves as the baseline for our framework. We further evaluate our method under two configurations: (Ours w/o Mem) — the feedback-driven model after adaptation training without memory loading; and (Ours w/ Mem) — the model evaluated with the reloadable memory bank.
@@
HAMT  [Chen_HAMT_2021] 40.12 46.25 38.23 5.74 3.97 7.72 55.54 34.06 56.09
NaviLLM  [Zheng_2024_CVPR] 49.92 60.25 43.23 4.65 3.01 10.25 57.15 39.29 56.44
DUET [Chen_DUET] 54.04 65.46 42.26 4.44 2.69 13.53 53.06 38.71 52.13
GR-DUET (baseline) [Hong_et_al_2025] 66.96 76.58 61.22 3.25 2.02 9.80 69.24 56.38 67.18
@@
Building upon GR-DUET, our proposed user-feedback-driven framework further enhances navigation performance across all evaluation metrics. The feedback-driven variant (Ours w/o Mem) achieves substantial improvement over GR-DUET, showing that incorporating user feedback significantly enhances both navigation success and path efficiency through explicit human-guided policy refinement. Enabling the reloadable memory bank (Ours w/ Mem) yields significant performance gains, demonstrating that previously explored environmental knowledge can be effectively leveraged to overcome the cold-start condition. Without loaded memory, the model begins with limited environmental context, leading to temporarily suboptimal performance that improves as memory is reconstructed during navigation. Overall, these findings suggest that user feedback acts as the primary driver of continual policy improvement, whereas the memory-bank warm-start mechanism offers complementary benefits by mitigating cold-start degradation across deployment cycles. A qualitative visualization in Fig.3 further supports these results: compared to the baseline GR-DUET (left), our user-feedback-adapted model (right) more accurately executes the instruction “walk past the couch and turn left, walk past the stairs and turn right, walk into the room and stop,” successfully reaching the target room. This visual evidence aligns with the consistent improvements across all quantitative metrics in Table I.
@@ IV-D Ablation Study
These results validate the complementary effects of exploration and feedback in our framework. Increasing the number of executed instructions enhances the agent’s spatial understanding, while additional feedback further refines its semantic grounding. Even under limited interaction (e.g., 100 feedback samples), the model achieves strong gains, demonstrating that our user-feedback-driven adaptation is both effective and sample-efficient.
@@ IV-F Continual and Hybrid Adaptation Results
TABLE V: Hybrid adaptation with five concurrent users. Full-metric comparison against GR-DUET baseline and continual adaptation (Adapt-3).
@@
In the hybrid setting, feedback from five users (100 samples each, 500 total) is merged into a single adaptation step. Despite using only one-third of the data and training duration of the continual setup, hybrid adaptation still yields notable improvements over GR-DUET—+4.47 SR, +6.00 SPL, and -0.53 m NE—with consistent gains across nDTW, SDTW, and CLS (Table V). This demonstrates that multi-user feedback aggregation can provide an efficient and scalable solution for real-world deployments where multiple users interact concurrently under limited training budgets.
Taken together, the two adaptation modes reflect distinct but complementary deployment scenarios—continual adaptation aligns with long-term, sequential user interactions, whereas hybrid adaptation represents concurrent multi-user adaptation within a single update cycle. Across both settings, our framework consistently yields substantial improvements over the GR-DUET baseline in all major metrics, demonstrating its robustness to diverse adaptation conditions and its capacity to effectively integrate user feedback.
@@ V Conclusion
In this work, we proposed a user-feedback-driven adaptation framework for Vision-and-Language Navigation (VLN), extending the General Scene Adaptation (GSA-VLN) paradigm beyond environment-only learning. Building upon GR-DUET, our framework systematically integrates explicit user feedback into the adaptation process and introduces a memory-bank-based warm-start mechanism for continual deployment. Through extensive experiments on the GSA-R2R benchmark, we demonstrated that: (1) user feedback provides highly effective supervision, substantially improving navigation success and path efficiency even with limited samples; (2) the memory-bank warm start mitigates cold-start degradation during redeployment by reusing environment-specific knowledge accumulated from prior sessions; and (3) both continual and hybrid adaptation modes enable consistent improvements over strong baselines, confirming the practicality of our approach under realistic long-term operation. Our results highlight the importance of leveraging real-world user interaction as a scalable and authentic source of supervision for embodied agents.

### v2

@@ III-C Stage 2: User-Feedback-Driven Trajectory Correction
Following the episode-level triggering condition in Sec. III-A (Eq. 67), the agent executes an instruction $I$ until termination at $T$, producing a trajectory
@@
with terminal viewpoint $v_{T}$. User feedback is queried only at this terminal step (i.e., $\mathrm{Tr}(T)=1$). Let $\hat{v}=\mathcal{F}(v_{T},v^{*})$ denote the supervision endpoint (Sec. III-A, Eq. 8), which is either the corrected goal $v^{*}$ (upon failure) or the confirmed stopping viewpoint $v_{T}$ (upon success). We then convert this sparse feedback to trajectory-level supervision by computing a feasible path on the current topological graph $G$:
$\tau^{+}=A^{*}(v_{1}\to\hat{v},G).$ (11)
Each corrected pair $(I,\tau^{+})$ is stored as a new adaptation sample.
To ensure data quality, we apply a length-based outlier rejection on the corrected trajectory $\tau^{+}$. Let $\lvert\tau^{+}\rvert$ denote the number of nodes (viewpoints) in $\tau^{+}$. Importantly, the thresholds are not chosen heuristically or tuned on downstream performance; instead, they are determined directly from the trajectory-length distribution of our corrected paths. Across 12,000 corrected trajectories constructed in Stage 2, the path length ranges from 1 to 14 with an average of 6.01. The distribution is highly concentrated: 94.5% of trajectories fall within 5–7 nodes (length 5: 24.2%, length 6: 34.8%, length 7: 35.5%). Therefore, we keep only trajectories with $5\leq\lvert\tau^{+}\rvert\leq 7$ (discarding $\lvert\tau^{+}\rvert<5$ or $\lvert\tau^{+}\rvert>7$), removing only 5.5% outliers, which are often caused by incomplete topology or corrections inconsistent with the discovered graph.
III-D Stage 3: Feedback-Based Adaptation and Memory-Bank Warm Start
@@ III-D1 Memory-bank Warm Start
After fine-tuning, the agent is redeployed in the same environment. Instead of reconstructing the graph and visual features from scratch, it reloads the previously stored memory $\mathcal{M}^{\prime}=(G,\mathcal{C},\mathcal{S})$, enabling immediate reuse of environment topology and cached representations. This reloadable memory extends the GR-DUET design by introducing persistent scene states that can be saved and restored across sessions, maintaining structural and perceptual continuity during redeployment.
@@ III-D3 Hybrid Adaptation
In practical deployments such as households, the same agent may serve multiple users in the same environment, and their feedback may accumulate within a deployment window. To handle such aggregated supervision efficiently, we introduce hybrid adaptation, which aggregates feedback from all users within a deployment window:
@@
where $\mathcal{D}_{u}$ represents feedback-corrected trajectories from user $u$. By jointly optimizing across multiple user datasets, hybrid adaptation captures diverse linguistic expressions and navigation styles while avoiding bias toward early interactions.
Consequently, the proposed framework supports two distinct update modes: sequential updates for continuous adaptation (Eq. 14) and batched aggregation for multi-user feedback (Eq. 15). This flexibility allows the agent to handle varying streams of supervision signals consistent with real-world deployment logistics.
@@ IV-A Experimental Setup
We conduct experiments on the GSA-R2R dataset [13], which extends the standard Room-to-Room (R2R) benchmark [2] to support scenario adaptation. In this work, we focus on the residential environments and evaluate on two instruction styles: Basic and User. This setting captures the most common real-world deployment scenarios, where navigation agents repeatedly operate in household layouts and encounter both neutral and user-specific language patterns.
@@
We adopt GR-DUET [13] as our primary baseline, since it represents the state of the art for GSA-VLN by explicitly maintaining a global navigation graph for adaptation. To evaluate our method against existing adaptation strategies, we include representative deployment-time adaptation (TTA) baselines built upon GR-DUET, such as entropy-minimization methods (TENT, FSTTA) and the feedback-aware ATENA. Finally, to further validate the generality of our approach, we also test other representative VLN methods under the GSA-VLN protocol, including DUET [8], HAMT  [7], and NaviLLM [40]. These methods were originally developed for the standard zero-shot VLN setting, and their performance in the adaptation scenario serves as an additional reference.
@@
All adaptation experiments share consistent optimization and hardware settings. We optimize the model using the AdamW optimizer with a learning rate of $1\times 10^{-5}$ and a batch size of 2. Each adaptation stage consists of 30,000 episodes. Training is conducted on a single NVIDIA RTX 4090 GPU for about 82 hours per stage.
@@ IV-B Evaluation Metrics
We evaluate navigation performance using standard metrics from the Vision-and-Language Navigation (VLN) literature, covering goal achievement, path efficiency, trajectory fidelity, and semantic alignment. Following the Room-to-Room (R2R) benchmark [2], we report nine widely used metrics: Success Rate (SR) [2], Oracle Success Rate (Oracle SR)[2], Success weighted by Path Length (SPL) [1], Navigation Error (NE) [2], Oracle Error / Oracle Navigation Error [16], Path Length [2], Normalized Dynamic Time Warping (nDTW) [15], Success weighted by nDTW (SDTW) [15], and Coverage weighted by Length Score (CLS) [16]. These metrics are adopted consistently across recent VLN benchmarks [2, 20, 16, 13] to ensure comparability with prior work.
@@ IV-C Main Results on the GSA-Basic Test Set
Table I presents the quantitative comparison of different VLN methods on the GSA-Basic test split. We include representative baselines HAMT [7], NaviLLM [40], DUET [8], and GR-DUET [13], where GR-DUET serves as the baseline for our framework.
We further include several deployment-time adaptation baselines on top of GR-DUET: entropy-minimization-based methods (TENT and FSTTA) and feedback-aware adaptation methods (ATENA). For all TTA baselines, we perform a hyperparameter sweep and report the best configuration under our evaluation protocol. We also evaluate our method under two configurations: the feedback-driven model after adaptation training without memory loading (Ours w/o Mem) and the model evaluated with the reloadable memory bank (Ours w/ Mem).
@@
HAMT  [7] 40.12 46.25 38.23 5.74 3.97 7.72 55.54 34.06 56.09
NaviLLM  [40] 49.92 60.25 43.23 4.65 3.01 10.25 57.15 39.29 56.44
DUET [8] 54.04 65.46 42.26 4.44 2.69 13.53 53.06 38.71 52.13
GR-DUET (baseline) [13] 66.96 76.58 61.22 3.25 2.02 9.80 69.24 56.38 67.18
GR-DUET+TENT [34] 67.46 76.96 61.65 3.23 2.01 9.69 69.47 56.84 67.30
GR-DUET+FSTTA [11] 67.33 77.17 61.37 3.21 1.99 9.87 69.12 56.50 67.17
GR-DUET+ATENA [18] 65.69 72.76 61.89 3.38 2.34 8.52 69.97 55.57 67.82
@@
Compared to GR-DUET, entropy-minimization-based TTA methods bring only marginal improvements (TENT: +0.50+0.50 SR / +0.43+0.43 SPL, FSTTA: +0.37+0.37 SR / +0.15+0.15 SPL), suggesting that self-supervised entropy signals are relatively weak in the GSA-VLN setting under distribution shift. ATENA presents a mixed trend. While it improves certain efficiency and fidelity metrics (e.g., SPL/nDTW/CLS) and achieves a shorter path length, it yields lower SR and Oracle SR in our setting. In contrast, our user-feedback-driven supervision provides consistent gains across almost all metrics.
Building upon GR-DUET, our proposed user-feedback-driven framework further enhances navigation performance across almost all evaluation metrics. The feedback-driven variant (Ours w/o Mem) achieves substantial improvement over GR-DUET, showing that incorporating user feedback significantly enhances both navigation success and path efficiency through explicit human-guided policy refinement. Enabling the reloadable memory bank (Ours w/ Mem) yields significant performance gains, demonstrating that previously explored environmental knowledge can be effectively leveraged to overcome the cold-start condition. Without loaded memory, the model begins with limited environmental context, leading to temporarily suboptimal performance that improves as memory is reconstructed during navigation. Overall, these findings suggest that user feedback acts as the primary driver of continual policy improvement, whereas the memory-bank warm-start mechanism offers complementary benefits by mitigating cold-start degradation across deployment cycles. A qualitative visualization in Fig. 3 further supports these results: compared to the baseline GR-DUET (left), our user-feedback-adapted model (right) more accurately executes the instruction “walk past the couch and turn left, walk past the stairs and turn right, walk into the room and stop,” successfully reaching the target room. This visual evidence aligns with the consistent improvements across all quantitative metrics in Table I.
@@ IV-D Ablation Study
We observe that environment exploration and user feedback play distinct but reinforcing roles. Increasing the number of executed instructions establishes a more comprehensive spatial representation, which in turn amplifies the effectiveness of semantic corrections provided by user feedback. Crucially, this combination enables robust adaptation even under limited supervision (e.g., 100 samples), demonstrating high sample efficiency.
@@ IV-E Effect of Executed Instruction Quantity
A concern in utilizing episode-level user feedback is whether accepting all trajectories that reach the correct target $v^{*}$, regardless of their intermediate path, introduces noise that hampers learning. To verify this, we conducted an additional analysis on the GSA-Basic split. We filtered the collected feedback samples to retain only those strictly matching the ground-truth shortest paths (reducing the dataset size from $\sim$11342 to 10,587 samples) and retrained the model. The results showed that the strictly filtered model achieved an SR of 76.33% and SPL of 71.14% at convergence, nearly identical to our proposed method (SR 76.33%, SPL 70.99%). This suggests that: (1) trajectories that successfully reach the target but deviate slightly from the oracle path still provide valid navigational signals; and (2) our feedback mechanism is robust to path variations, effectively leveraging episode-level supervision without requiring strict step-by-step alignment, which is often unavailable in real-world scenarios.
@@ IV-F Continual and Hybrid Adaptation Results
TABLE V: Hybrid adaptation with five users on the User (mixed) split. Full-metric comparison against GR-DUET under the same evaluation protocol.
@@
TABLE VI: Training cost and optional memory-bank storage overhead.
Mode #Stages Train time Mem (MB) (w/o / w/)
Hybrid (User) 1 $\sim$82 h 0 / 16.32
Continual (Basic+User) 3 $\sim$246 h 0 / $\sim$35
@@
In the hybrid setting, feedback from five users (100 samples each, 500 total) is merged into a single adaptation step. Despite using only one-third of the data and training duration of the continual setup, hybrid adaptation still yields notable improvements over GR-DUET—+9.91 SR, +11.14 SPL, and -0.79 m NE—with consistent gains across nDTW, SDTW, and CLS (Table V). This demonstrates that multi-user feedback aggregation can provide an efficient and scalable solution for real-world deployments where multiple users provide feedback over a deployment window under limited training budgets.
As highlighted by the experimental results, our framework introduces additional components to enable adaptation. To assess the practicality of our method, we analyze the training cost and memory consumption in Table VI. Under identical optimization and hardware settings, each adaptation stage requires approximately 82 hours of training; therefore, Hybrid adaptation performs a single aggregated update within a deployment window and costs 82 hours in total, whereas Continual adaptation updates the agent sequentially across stages and scales roughly linearly with the number of stages (e.g., three stages cost 246 hours). In our reported Continual/Hybrid experiments, we did not enable the persistent memory-bank warm start (save/load across stages), and thus there is no extra persistent storage overhead (0 MB). The memory-bank warm start is an optional deployment feature, and its storage depends on the set of scans/environments that the agent needs to cache. In our setup, the Basic split covers 24 scans (18.62 MB) while the User split covers 21 scans (16.32 MB). Therefore, Hybrid adaptation (User-only) would require 16.32 MB if warm start is enabled. For Continual adaptation, we run one stage on Basic scans and two stages on User scans; Since the two User stages share the same scan set but differ in viewpoint coverage, enabling warm start would only require storing the scan set once, totaling approximately 35 MB.
@@ V Conclusion
In this work, we introduced a user-feedback-driven adaptation paradigm for Vision-and-Language Navigation (VLN), fundamentally shifting the reliance from unstable environment-driven self-supervision to intent-aligned human guidance. By leveraging a topology-aware trajectory construction pipeline, our framework effectively “lifts” sparse, episode-level corrections into dense, grounded supervision, enabling sample-efficient policy refinement without requiring dense step-by-step demonstrations. We further demonstrated that a persistent, reloadable memory bank significantly mitigates cold-start degradation during redeployment by preserving acquired structural and perceptual knowledge. Extensive experiments on the GSA-R2R benchmark confirm that our approach consistently outperforms environment-driven baselines on the standard Basic split, while effectively handling diverse instruction styles in both continual and hybrid adaptation settings. This validates that our proposed user feedback mechanism serves as a robust and scalable supervision signal for realistic deployment scenarios. Although our experiments are conducted in static benchmark environments, the proposed feedback abstraction does not assume scene stationarity. In dynamic settings, the same (success, goal-correction) interface remains valid, while the trajectory lifting step can be performed on an updated topological map.